The width of a rectangle is 5 ft less than the length. The area of the rectangle is 300 ft. How do I find the length and width of the rectangle?
Answer by Cromlix(4381)

Length = x
Width = x - 5
Area = length x width
Area = 300ft^2
x(x - 5) = 300
x^2 - 5x - 300 = 0
(x + 15)(x - 20) = 0
x + 15 = 0
x = -15 (no answer, as negative)
x - 20 = 0
x = 20
Length = 20 ft
Width = 15 ft.
Hope this helps.
:-)